OpenStudy (anonymous):

Joshua opened a savings account 11 years ago. The interest rate was 2.9% compounded quarterly. If his balance is currently $1,437.34, how much did he originally deposit into the account?

OpenStudy (anonymous):

$1,045.97 $1,089.72 $1,265.39 $1,396.83

OpenStudy (anonymous):

wait one sec i'll get it

OpenStudy (anonymous):

Okay. Thank you!

OpenStudy (anonymous):

i'm not completely sure if this is correct

OpenStudy (anonymous):

first i would take each number and multiply by 2.9

OpenStudy (anonymous):

i would say that the answer is 1,265.39

OpenStudy (anonymous):

Well, I did 1437.34=p(1+0.00725)^11 1437.34=p(1.00725)^11 1437.34=p(1.08270474) then divide 1.08270474 from both sides and I got p=1327.55 and that isn't one of the answers.

OpenStudy (anonymous):

I got 0.00725 by dividing 0.029 by 4 because it is compounded quarterly.